A factory is discharging wastewater containing 200 mg/L BOD into a river. The river has a discharge of 10 m3/s and the wastewater discharge is 2 m3/s. If the upstream of this river has a BOD of 10 mg/L. Please compute BOD after wastewater is completely mixed with river (any BOD degradation is not considered)
Answer #1

The discharge is =10 cubic metre/sec

So 10×1000 lites =1000 litres

Upstream BOD = 10mg/litre

Now after waste water total discharge will be

10+2 =12 cubic metre/sec

= 12000 litre

Total Factory discharge = 2 cubic metres/sec

= 2000 litre

BOD = 200 mg/litre

So total BOD after Water water Mixing

= 10000×10 + 2000×200/12000 litres

100,000+400,000mg /12000 litre

= 500,000/12000 = 41.66 mg/litre

Please upvote the answer.....